Determine the all of the voltages and the currents, including the total current, of this circuit. Also draw an equivalent circuit. 125 V 20 ? 100 ? 50?
Answer #1

As the three resistors connected parallel to the battery potential deiffrence across them must be same.

potential diffrence across all the three resistors = 125 V

current through 20 ohm = 125/20 = 6.25 A

current through 100 ohm = 125/100 = 1.25 A

current through 20 ohm = 125/50 = 2.50 A


1/Req = 1/20 + 1/100 + 1/50

Req = 12.5 ohms

For equivalent circuit, replace the three resistors with a single resistor with resistance 12.5 ohms